# LongGenbench: Benchmarking Long-Form Generation in Long Context LLMs

### v6

@@ 2.1 Task Definition
Evaluating the quality of super-long-form generation  presents a unique set of challenges due to the inherent complexity of long texts. Traditional human evaluation methods, while precise, are expensive and not scalable. Although using large language models for assessment is feasible, their lack of interpretability often hampers their utility. Thus, we focus on the ”instruction-following” task in super-long-form generation, where the most must include specific details in the generated text. This task reflects real-world scenarios that require a high degree of attention to detail over extended sequences, such as technical documentation or detailed design proposals. In this study, we define a task type termed Strictly Sequential Tasks, which involves the sequential completion of subtasks $\mathbf{T}=\{T_{1},T_{2},T_{3},\ldots,T_{n}\}$, where each subtask is responsible for generating a specific volume of text. For instance, an LLM might be tasked with designing a 100-floor skyscraper, specifying the content and purpose of each floor.
@@ 2.2 Four Distinct Scenario Setups
These scenarios were carefully chosen to reflect both creative and technical long-form generation tasks. They offer a diverse set of challenges by including temporal tasks (e.g., Diary Writing) that require maintaining consistent information over time and spatial tasks (e.g., Urban Planning) that test the model’s ability to handle spatial relationships and detailed designs. These scenarios mirror real-world applications, from planning documents to creative writing, and thus provide a comprehensive evaluation of long-context LLMs. Table 2 offers comprehensive descriptions for each scenario, with each designed around a unique template to generate 100 different task instructions [^2] Examples of Task Instructions for each scenario are provided in Appendix A..
@@ 2.3 Specific Task Instruction
Completion Rate=Number of Completed SubtasksTotal Number of $\text{Completion Rate}=\frac{\text{Number of Completed Subtasks}}{\text{Total Number of Subtasks}}\times 100\%$
@@
STIC-1 is primarily concerned with the completion rate of instructions that result in sub-scenarios, focusing on whether instructions are correctly executed. In contrast, STIC-2 assesses the overall completion of the specific instruction task, including the presence of sub-scenarios and their completion status.
@@
Our evaluation process follows a structured pipeline: First, we use a long-context LLM to complete the task instruction $T$, generating an answer $A$, which is then divided into sub-tasks as $A=\{A_{1},A_{2},\ldots,A_{n}\}$. Next, based on the specific instructions in the check_set, we identify the relevant sub-tasks within $A$. Finally, we evaluate each sub-task by $\text{eval}(A_{i},T_{i})$ to compute the final completion score, as detailed in Algorithm 1. This pipeline ensures that the evaluation is both systematic and comprehensive, assessing the model’s performance across different instruction settings and levels of complexity. While LongGenBench primarily evaluates the model’s ability to follow detailed instructions, future work could expand the benchmark to include more open-ended tasks that assess creativity and logical reasoning. This would provide a broader evaluation of a model’s capabilities in generating coherent, engaging, and logically sound long-form text.
@@ 3.1 Experimental Setup
We selected ten long-context large language models (LLMs), comprising eight open-source and two closed-source models. These models range in size from 7B to 72B parameters, with one featuring a Mixture of Experts (MoE) architecture. The claimed context lengths of these models vary from 32K to 128K tokens [^4] Detailed specifications of these models are provided in Appendix C.. These models were selected to represent a diverse array of architectures, covering both Mixture of Experts and standard transformer designs, as well as a range of parameter sizes. This diversity ensures a comprehensive evaluation of their ability to handle long-context tasks.
@@ 3.2 Main Result
Table 3: Long-form generation Performance (%) of selected models evaluated at length from 16k and 32k. The weighted average score (wAvg) is the product of CR and STIC-2, used to represent the model’s final performance at the given task length. Note that the GPT-4-32K is currently closed for use, and the longest versions that can be used are the GPT-4o and GPT-4o-mini 16K output limitation.
@@
Results.
The results of the long-form text generation tasks for both 16K and 32K tokens are summarized in Table 3. The models exhibited significant performance variations, highlighting differences in architecture, parameter size, and training data.
@@ Results.
A common failure mode observed across multiple models was the tendency to forget or misinterpret instructions as the sequence length increased. For example, in the Skyscraper Design task, some models correctly described the initial few floors but deviated from the original plan as the task progressed, particularly in the 32K token setting. This highlights the memory retention issue in long-context models, which often leads to a loss of coherence and adherence to task instructions. Examples of these failures, where models struggled to follow prompt instructions over extended outputs, are provided in Appendix D.
@@ 6 Conclusion
Appendix A PROMPT TEMPLATES FOR FOUR TASK SCENARIOS
@@
Appendix B evaluation prompts
We provide the prompts used for evaluation and the function $\text{eval}(A_{i},T_{i})$ in Algorithm 1, where Corresponds to the output part of Specific Task Instruction represents $A_{i}$, and Specific Task Instruction represents $T_{i}$. We present both a ”yes” example and a ”no” example for the few-shot evaluation.
Evaluation Prompts Example 1 XXXX Answer: yes Example 2 XXXX Answer: no Context: ”Corresponds to the output part of Specific Task Instruction” Instructions: Does this context include Specific Task Instruction? Please answer with ’yes’ or ’no’ only.”
Appendix C Models
@@
Appendix D Error Analysis

### v7

@@ 2.1 Task Definition
Evaluating the quality of super-long-form generation  presents a unique set of challenges due to the inherent complexity of long texts. Traditional human evaluation methods, while precise, are expensive and not scalable. Although using large language models for assessment is feasible, their lack of interpretability often hampers their utility. Thus, we focus on the ”instruction-following” task in super-long-form generation, where the most must include specific details in the generated text. This task reflects real-world scenarios that require a high degree of attention to detail over extended sequences, such as technical documentation or detailed design proposals. In this study, we define a task type termed Strictly Sequential Tasks, which involves the sequential completion of subtasks $\mathbf{T}=\{T_{1},T_{2},T_{3},\ldots,T_{n}\}$ [^2] In Appendix A, there is a detailed description of the definitions of mathematical symbols., where each subtask is responsible for generating a specific volume of text. For instance, an LLM might be tasked with designing a 100-floor skyscraper, specifying the content and purpose of each floor.
@@ 2.2 Four Distinct Scenario Setups
These scenarios were carefully chosen to reflect both creative and technical long-form generation tasks. They offer a diverse set of challenges by including temporal tasks (e.g., Diary Writing) that require maintaining consistent information over time and spatial tasks (e.g., Urban Planning) that test the model’s ability to handle spatial relationships and detailed designs. These scenarios mirror real-world applications, from planning documents to creative writing, and thus provide a comprehensive evaluation of long-context LLMs. Table 2 offers comprehensive descriptions for each scenario, with each designed around a unique template to generate 100 different task instructions [^3] Examples of Task Instructions for each scenario are provided in Appendix B..
@@ 2.3 Specific Task Instruction
[…]
@@
STIC-1 is primarily concerned with the completion rate of instructions that result in sub-scenarios, focusing on whether instructions are correctly executed. In contrast, STIC-2 assesses the overall completion of the specific instruction task, including the presence of sub-scenarios and their completion status [^5] In Appendix E, we provide a detailed explanation of STIC-1 and STIC-2, along with a case study analysis..
@@
Our evaluation process follows a structured pipeline: First, we use a long-context LLM to complete the task instruction $T$, generating an answer $A$, which is then divided into sub-tasks as $A=\{A_{1},A_{2},\ldots,A_{n}\}$. Next, based on the specific instructions in the check_set, we identify the relevant sub-tasks within $A$. Finally, we evaluate each sub-task by $\text{eval}(A_{i},T_{i})$ to compute the final completion score, as detailed in Algorithm 1. This pipeline ensures that the evaluation is both systematic and comprehensive, assessing the model’s performance across different instruction settings and levels of complexity  [^6] In Appendix C, there is a detailed evaluations pipeline and example.. While LongGenBench primarily evaluates the model’s ability to follow detailed instructions, future work could expand the benchmark to include more open-ended tasks that assess creativity and logical reasoning. This would provide a broader evaluation of a model’s capabilities in generating coherent, engaging, and logically sound long-form text.
@@ 3.1 Experimental Setup
We selected ten long-context large language models (LLMs), comprising eight open-source and two closed-source models. These models range in size from 7B to 72B parameters, with one featuring a Mixture of Experts (MoE) architecture. The claimed context lengths of these models vary from 32K to 128K tokens [^7] Detailed specifications of these models are provided in Appendix D.. These models were selected to represent a diverse array of architectures, covering both Mixture of Experts and standard transformer designs, as well as a range of parameter sizes. This diversity ensures a comprehensive evaluation of their ability to handle long-context tasks.
@@ 3.2 Main Result
Table 3: Long-form generation Performance of selected models evaluated at length from 16k and 32k. The weighted average score (wAvg) is the product of CR and STIC-2, used to represent the model’s final performance at the given task length. Note that the GPT-4-32K is currently closed for use, and the longest versions that can be used are the GPT-4o and GPT-4o-mini 16K output limitation.
@@
The results of the long-form text generation tasks for both Short-version (16K) and Long-version (32K) tokens are summarized in Table 3.
@@
A common failure mode observed across multiple models was the tendency to forget or misinterpret instructions as the sequence length increased. For example, in the Skyscraper Design task, some models correctly described the initial few floors but deviated from the original plan as the task progressed, particularly in the 32K token setting. This highlights the memory retention issue in long-context models, which often leads to a loss of coherence and adherence to task instructions. Examples of failures where models struggled to follow instructions are provided in Appendix  F.
@@ 6 Conclusion
Appendix A Symbol Definitions and Descriptions
This table 4 presents definitions and descriptions for various symbols used in task-related contexts, providing an overview of the key terminologies and their roles.
Symbol Definition Description
$T$ Main Task The primary goal or task to be completed, such as designing a skyscraper or writing a diary.
[…]
Table 4: Symbol Definitions and Descriptions
Appendix B PROMPT TEMPLATES FOR FOUR TASK SCENARIOS
@@
Appendix C Evaluation pipeline
The evaluation pipeline is designed to systematically assess the ability of long-context language models (LLMs) to follow specific, complex instructions. The process can be summarized in three key steps:
Step 1. Generation of Outputs from the Long-context LLM
Given an input task ($T$) that describes a set of instructions, we prompt the LLM to generate detailed outputs. The output ($A$) comprises a list of descriptions, represented as:
$A=\{A_{1},A_{2},\dots,A_{n}\}$
Example: Given the prompt (ref Appendix B) Construct a skyscraper with 100 floors. The floor assignments are detailed as follows: Specific floor requirement: Designate Floor 11 for a small art gallery. Range floor requirement: Allocate Floors 32 to 39 for corporate headquarters of a major company.
The LLM generates a response describing each floor in detail, such as:
Answer: Floor 1: …Lobby … Floor 11: …Small art gallery … Floor 32: …Corporate headquarters … Floor $n$: …
Step 2. Extracting and Matching Relevant Floor Assignments (Check Set)
From the initial input (“T”), we create a check set containing specific floor assignments to verify if the LLM correctly follows the instructions.
For the example above, the check set includes:
Check Set: Floor 11: Small art gallery Floor 32: Corporate headquarters Floor 33: Corporate headquarters
We then extract the relevant parts of the LLM output (“A”) that correspond to the floor assignments described in the check set.
Step 3. Evaluation Using Llama 3.1-8B instruction Model
For each extracted pair, we use the Llama 3.1-8B model to evaluate whether the output (“$A_{i}$”) for a given task segment (“$T_{si}$”) has correctly fulfilled the specified instruction.
This evaluation task is framed as a simple binary classification problem, which aims to determine if the specific instruction was fulfilled (“yes” or “no”). The prompt used for this evaluation is as follows:
Evaluation Prompts Example 1: XXXX Answer: Analysis + #*# Yes Example 2: XXXX Answer: Analysis + #*# No Context: Long-context model output: ”Floor 11: …small art gallery …” Instructions: Does this context include ‘small art gallery’? Answer: Please refer to the above example, provide your analysis, and respond with either #*# Yes or #*# No.
Notably, this binary evaluation is straightforward. We manually labeled 300 data points, and the model’s output matched human evaluations for all cases.
By using this process, we transform the evaluation of long-context text generation into multiple evaluations of smaller segments. This enables systematic and thorough verification of how well the LLM follows the instructions for each specific task (as detailed in the check set).
Appendix D Models
@@
Appendix E Explanation of Metrics
E.1 STIC-1 and STIC-2
We appreciate the feedback and have provided an example using the results from Table 3 of our experiments (specifically comparing LLaMA3.1-8B and Qwen2 under the short-version setting).
As shown in Table 6, Qwen2’s STIC-1 score is higher than that of LLaMA3.1-8B, while its STIC-2 score is lower. This difference can be attributed to the Completion Rate (CR) of each model. Qwen2 has a significantly lower CR compared to LLaMA3.1-8B. Specifically, Qwen2 typically achieves around 60% completion for tasks (e.g., designing a 100-story skyscraper but stopping at roughly 60 stories). On the other hand, LLaMA3.1-8B generally completes around 93 layers (93% completion).
In the case of STIC-1, we are evaluating the correctness of the output based on the number of layers that are actually generated. Qwen2 demonstrates a higher completion rate when the denominator consists of the 60 layers it has output (compared to LLaMA3.1-8B, which has a denominator of 93 layers).
For STIC-2, however, we consider the entirety of the expected output. Since Qwen2 lacks the remaining 30 layers, the STIC-2 score is lower when the denominator becomes the entire requirement (as the missing output significantly affects its score).
As mentioned in our paper, STIC-2 is designed to take into account a more comprehensive perspective on output completeness. We are considering simplifying our metrics by using only STIC-2, as it may be easier to understand and provide a more holistic evaluation.
Table 6: Comparison of STIC-1 and STIC-2 Scores
Model Length CR STIC-1 STIC-2
LLaMA3.1-8B 128K 93.5% 23.4% 22.0%
Qwen2-7B 128K 60.0% 27.9% 16.1%
E.2 Example explanation
To further illustrate the concepts, we have constructed a 3-level building for illustration [^8] This example is adapted from ICLR reviewer bHUs. We are deeply grateful for their insightful comment and in-depth discussion, which significantly improved the clarity of this paper.:
  • Consider a 3-level building with the following constraints:
    • $T_{S_{1}}$: “Floor 1 must have a coffee shop.”
    • $T_{S\_2}$: “Floor 1 must have a reception desk.”
    • $T_{P}$: $\{T_{P_{1}},T_{P_{2}},T_{P_{3}}\}$, where each $T_{P_{i}}$ means “Floor $i$ must have a washroom.”
The model generates the following output:
“floor1: coffee shop, washroom; floor2: washroom.”
In this scenario, the check_set is $\{T_{S_{1}},T_{S_{2}},T_{P_{1}},T_{P_{2}},T_{P_{3}}\}$. Note that $T_{P}$ applies to all three floors, requiring separate evaluation for each $TP_{i}$.
With the current model output, the completion rate (CR) for the main task is $2/3$. Although the task requires outputs for three floors, the model only provided outputs for two floors.
For STIC-1, we consider how accurately the model has outputted information at the floor level. Since the model output only contains two floors, we evaluate the constraints for these two floors to determine if they are fully met. For these two floors, the constraints are $T_{S_{1}}$, $T_{S_{2}}$, $T_{P_{1}}$, $T_{P_{2}}$, totaling 4 constraints. The model has correctly fulfilled 3 out of these 4 requirements, resulting in STIC-1 of $3/4$.
For STIC-2, we evaluate the entire check_set, which consists of $T_{S_{1}}$, $T_{S_{2}}$, $T_{P_{1}}$, $T_{P_{2}}$, $T_{P_{3}}$. The model has fulfilled 3 out of these 5 requirements, so STIC-2 equals $3/5$.
The distinction between STIC-1 and STIC-2 allows identification of the specific reasons for any drop in performance. It helps determine whether the issue lies in the model’s inability to follow instructions for a given output or whether it lacks a complete output in the first place. For example, in the case of a lower STIC-2, the low score may be due to incorrect outputs for some floors or due to incomplete outputs for the floors. In such cases, CR and STIC-1 can be used together to further evaluate and make informed judgments.
Appendix F Error Analysis
@@
Appendix G Different Prompt Format Compare
The two prompt formats differ primarily in the structure and arrangement of the instructions within the prompt. In Prompt - 1, the order follows a sequence of Single Instruction (SI), Range Instruction (RI), and Periodic Instruction (PI). Conversely, in Prompt - 2, this order is altered by swapping the positions of SI, RI, and PI. Additionally, the Generate prompt, which is a critical component of the task, was rewritten in Prompt - 2 by GPT-4.
From the table, it is evident that different prompt formats influence the performance metrics of the models, such as CR, STIC-2, length, and wAvg. For instance, in Prompt - 1, the Mistral-7B-Instruct model achieves the highest CR (81.8) and STIC-2 (17.44%), while LongWriter-llama3.1-8b lags behind with a CR of 46.0 and STIC-2 of 9.83%. Similarly, under Prompt - 2, the same trend is observed: Mistral-7B-Instruct maintains its lead with a CR of 62.3 and STIC-2 of 16.29%, while LongWriter-llama3.1-8b again ranks lowest with a CR of 24.3 and STIC-2 of 8.35%.
Although the prompt format does affect the absolute values of these metrics (e.g., all models show reduced CR under Prompt - 2 compared to Prompt - 1), the relative rankings remain unchanged. This consistency suggests that while prompt design impacts performance, it does not alter the comparative effectiveness of the models.
Prompt Format Model CR STIC-2 Length (word) wAvg Rank
Prompt - 1 LongWriter-llama3.1-8b 46.0 9.83% 11036 4.5 3
Qwen2-7B-Instruct 60.0 16.13% 5138 9.7 2
Mistral-7B-Instruct-v0.2 81.8 17.44% 7296 14.3 1
Prompt - 2 LongWriter-llama3.1-8b 24.3 8.35% 6189 2.0 3
Qwen2-7B-Instruct 57.3 16.34% 4334 9.4 2
Mistral-7B-Instruct-v0.2 62.3 16.29% 4750 10.2 1
Table 7: Model comparison with different prompt formats
Appendix H Explanation of Unit Differences: Tokens vs. Words
In this work, the term “16K/32K” refers to the required number of tokens for the model output, adhering to the standard conventions when discussing model context lengths. However, for evaluating the actual generated output, we employed word count as the measurement unit. This distinction was made for the following key reasons:
Variability in Token-to-Word Conversion: Different tokenizers vary in how they convert tokens into words, typically resulting in an average ratio of about 1.5 tokens per word. Therefore, the actual word count of the output is usually approximately two-thirds of the target token length. This variability makes word count a more consistent measure for analyzing content.
Emphasis on Content Quality: Our primary focus was on evaluating the quality and completeness of the generated content. Word count provides a more straightforward perspective on the substance of the output, which is crucial for content assessment.